use induction to prove (1*2)+(2*3)+(3*4)+...+(n-1)*n= n(n-1)(n+1)/3 for all n >= 2.
Answer #1

Using induction proof is as follows

To prove, 12+2 3+3 4 + ... + (n-1). A = n(n-1) (n+1) 9 ; n22 3 For n=2, L.H.S = 1x2 =2 R.H.S= 2(2-1)(2+1) - 2 3 So relation i